OpenStudy (izzyrawrz):

Assuming that all volume measurements occur at the same temperature and pressure, how many liters of water vapor will be produced in the reaction? Balanced Equation: C₃H₈ + 5 O₂ -> 3 CO₂ + 4 H₂O

OpenStudy (photon336):

similar idea in the other question

OpenStudy (photon336):

at standard temperature and pressure STP, 1 mole of every gas occupies 22.4L

OpenStudy (rushwr):

Do u still need help? @izzyrawrz

OpenStudy (izzyrawrz):

Yes please @Rushwr

OpenStudy (rushwr):

yeah so at STP every gas occupies a volume of 22.4 L So here water vapour is considered as a gas. In the reactant side you have 4 moles of Water vapour. If 1 mole occupies 22.4L , 4 moles will occupy 22.4 * 4 will be the volume occupies by 4 moles of water vapor

OpenStudy (rushwr):

So if 1 moles occupies 22.4 L Then 4 moles occupy 22.4 *4 = 89.6L

OpenStudy (izzyrawrz):

Thanks can you help me verify this one Assuming that all volume measurements occur at the same temperature and pressure, how many liters of oxygen will be required to completely burn 0.700 L of propane gas? I came up with 4 L of oxygen is this correct

OpenStudy (izzyrawrz):

Balanced equation is the same one as the first question

OpenStudy (rushwr):

oops sorry didn't see this one coming mm wait let me check !

OpenStudy (rushwr):

nop nop not 4l

OpenStudy (rushwr):

we know mole/stoicheometric ratio = volume ratio \[\frac{ volume of C3H8 }{ Volume of O2 }= \frac{ Moles ofC3H8 }{ Moles of O2 }\]

OpenStudy (rushwr):

\[\frac{ 0.7 L }{ Volume of O2 } = \frac{ 1 }{ 5}\]

OpenStudy (izzyrawrz):

So then it would be 3.5 L of oxygen?

OpenStudy (rushwr):

\[Volume of O2 = \frac{ 0.7 * 5 }{ 1 } = 3.5 L\]

OpenStudy (izzyrawrz):

Oh I know what I did wrong the first time...I rounded :l

OpenStudy (rushwr):

oh lol okai ! :)

OpenStudy (izzyrawrz):

So thank you for your time :)

OpenStudy (rushwr):

no problem :)